(1) Calculate ΔE, if q= 0.767 kJ and w= -870 J .

(2)Calculate ΔE, if a system releases 66.7 kJ of heat to its surroundings while the surroundings do 44.0 kJ of work on the system
1
Expert's answer
2016-10-15T15:09:04-0400
ΔE = q + w
1)ΔE = 0.767*10^3 + (-870) = -103 (J);
2)ΔE = -66.7 + 44.0 = -22.7 (kJ) = -22.7*10^3 (J)
Though the system loses heat, a value of 'q' was taken with minus.